OpenStudy (anonymous):

How do i find the asymptotes? y=x+3/x+2

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (johnweldon1993):

vertical asymptote = make the denominator = 0

OpenStudy (johnweldon1993):

horizontal asymptote = set x to infinity and see what the equation comes to

OpenStudy (johnweldon1993):

you have both a horizontal and vertical asymptote so there will be no slant asymptote

OpenStudy (tkhunny):

As it is written, "y=x+3/x+2", you have a vertical asymptote at x = 0 and an oblique asymptote at y = x + 2. If you recall your Order of Operations, and write y = (x+3)/(x+2), that is quite a different problem.

OpenStudy (comm.dan):

There will be a x and a y asymptote. They will both be drawn using dashed lines.